# Segregation at prior austenite grain boundaries: the competition between boron and hydrogen

### v1

@@ 3.1 Experimental details
The incorporation of hydrogen into $LC$ and $B-LC$ has been performed electrochemically, using a three-electrode system on 10$\times$10$\times$1 mm3 samples. The procedure consisted of imposing a current of -25 mA.cm-2 for 1 h in a 0.1 M $H_{2}SO_{4}$ aqueous solution with 0.3 wtof NH4SCN. After charging, the oxide layers were removed from the sample using 600 SiC grinding paper before the TDS measurements. For all conditions, the time between the end of the electrochemical charging and the beginning of the temperature ramping is lower than 5 min, which is performed in an infrared furnace from 298 K to 1173 K with a mass spectrometer using the G8 GALILEO apparatus from Bruker®. The heating rate $\varphi$ varied between 8 $K.min^{-1}$ and 32 $K.min^{-1}$ to locate where hydrogen is trapped in the microstructure.
@@ 3.2 Results
Figure 4: TDS spectra of LC and B-LC heated for different temperature ramping at a) 8 K.min-1, b) 12 K.min-1, c) 16 K.min-1, d) 20 K.min-1, e) 24 K.min-1, and f) 32 K.min-1.
@@
Fig. 5.a presents the Kissinger plot of the different peaks shown in fig. 4 for LC and B-LC steels. The first peak from the TDS spectra of fig. 4 should correspond to mobile hydrogen in the lattice (from the slope of the blue curve). The energy obtained is -0.16 eV and is lower than the theoretically expected activation barrier for diffusion of hydrogen in martensitic steel or $\alpha$-iron, which is between -0.04 eV and -0.1 eV [72, 39, 70]. The main reason is that the determination of this energy comes from the temperature at which the hydrogen desorption rate $\phi$ is the highest. When the heating rate ($\varphi$) is low, this temperature should come around or below room temperature [44], the starting temperature from our experiments. The second peak (red curves) has been attributed to hydrogen trapped at GBs in previous work [38, 44]. Our results show a fourth peak which pertains to hydrogen desorbing from PAGBs. Consequently, this second peak should mostly be related to the desorption of hydrogen trapped in grain boundary from martensite transformation, which follows the KS OR, and have the same parent austenite grain. The trapping energy deduced from our experiment is -0.26 eV for hydrogen segregating into this type of trapping site (deduced from the slope of the red curve). This energy is in the range of the ones reported in the work of Galindo et al. [40], but outside the range given in Frappart et al. [39, 69] (fig. 5.b). The slope of the green curve gives the average trapping energy of -0.34 eV, and should represent mostly hydrogen in dislocations according to the literature [38, 44]. This value is in the range of the energies given in previous studies [39, 45, 40]. The last peak, observed only for the boron-free steel (orange curve in fig. 5.b) should be hydrogen in PAGBs with a trapping energy of -0.37 eV. This value is close to one reported in the work of Bhadeshia [45] but higher than the one reported in the work of Frappart et al. [39].
@@
In addition, hydrogen could also be trapped in the 2% of retained austenite observed in LC and B-LC steels and in or at carbides [45, 40, 67], precipitates observed by ECCI and synchrotron X-ray measurements. When hydrogen is trapped in austenite, a peak from the TDS spectra above 573 K should be obtained [67], which is not the case, neither for LC nor for B-LC. From the synchrotron experiments, a lattice strain of -0.275% is observed, which is equivalent to a compressive stress of 1.2 GPa. One possible explanation is that the austenite, which is too compressed to transform into martensite, is becoming a repulsive site for hydrogen to segregate. Such effect is observed in the case of edge dislocations for BCC and FCC metals, where the stress field of the compressed area induces a negative binding energy for hydrogen, which favors its segregation into the tensile area of the defect [73, 74]. Additionally, the carbide concentration is too low to observe an apparent peak by TDS measurements. A deconvolution of the TDS spectra considering these defects could be done near peaks 2 and 3, but might lead to an inaccurate overfit. Another peak has been found for hydrogen in microvoids in previous works, which has an amplitude significantly lower than the other peaks [38, 44]. However, no additional peak has been observed in either of the two steels. One possible explanation is that hydrogen recombines in these microvoids, making it impossible for hydrogen to desorb at the temperature range studied in this work [44].
@@ 4.1 Computational details
Ab initio calculations have been performed to understand the interaction between hydrogen and boron at grain boundaries by determining the interaction energies between solutes and the structural defect. The structural defect investigated is a $\alpha$-$\Sigma$5(210) grain boundary, which can also be written as $\alpha$-39°[001](210) GB. It is a coherent site lattice grain boundary that does not follow the KS OR, thus it should be representative of one type of PAGBs than a boundary from martensitic transformations. The calculations have been carried out using the Vienna Ab initio Simulation Package (VASP) code [75, 76, 77] in the generalized gradient approximation (GGA) using the PBE functional [78]. A plane-wave cut-off of 500 eV has been taken for all calculations. The convergence tolerance of the atomic force is 0.01 eV/Å and the Brillouin-zone integration was made using Methfessel–Paxton smearing [79]. The magnetic state of $\alpha$-Fe is ferromagnetic and a $4\times 4\times 4$ supercell containing 128 Fe atoms has been used for the bulk calculations. For the reciprocal space integration, a $\Gamma$-centered Monkhorst-Pack [80] k-point sampling of $8\times 8\times 8$ has been used for bulk $\alpha$-Fe calculations. The lattice constant and magnetic moments of the ferromagnetic $\alpha$-Fe were computed as 2.834 Å and 2.2 $\mu_{\rm B}$/atom, respectively, in agreement with the previous studies [81]. The supercell containing the $\alpha$-$\Sigma$5(210)[001] grain boundary consisted of 80 Fe-atoms. For the reciprocal space integration, a $\Gamma$-centered Monkhorst-Pack [80] k-point sampling of $5\times 10\times 16$ has been used to ensure convergence. The solute atoms H, C, and B have firstly been incorporated in different locations of the defect (in interstitial sites $I_{1}$ and $I_{2}$), represented in fig. 6. Then, additional calculations have been performed to describe the interaction between a boron-doped GB with an additional solute atom. In this situation, boron has been incorporated in $I_{1}$ and a second solute atom (hydrogen and boron) has been incorporated in interstitial site $I_{2}$ or in interstitial sites in the vicinity of the GB (in octahedral and tetrahedral sites).
@@ 4.2 Results
The interaction energy between solutes and boron-doped GB ($E^{\rm inter}_{\rm X-B-GB}$) is presented in fig. 7.b. In this configuration, additional boron is attracted to the GB with interaction energies lower than -0.5 eV in $O_{2}$. This interaction energy suggests that boron can be accumulated in GBs, which facilitates the formation of borides when the nucleation barrier is low enough [29, 30]. In contrast, hydrogen is repelled from the GB in this configuration when inserted in interstitial sites. For these calculations, when hydrogen atoms are initially inserted in octahedral sites, they are moving to their closest stable tetrahedral site. When a second boron atom is initially inserted in the tetrahedral site, it moves towards its closest stable octahedral site.
@@ 5 Hydrogen concentration in steels at thermodynamic equilibrium
The hydrogen permeation study has been conducted with a Kelvin probe-based detection method developed in-house [87, 88]. On the side of the sample designated for hydrogen detection, a dry nitrogen atmosphere is maintained, and a 100 nm Pd coating was applied using a Leybold Univex 450 physical vapor deposition apparatus. The hydrogen-entry side of the sample has been subjected to the same galvanostatic condition as samples for TDS measurements. Time lags of 11.6 min, 15.1 and 17.6 min are obtained for LC and 6.8 min, 8.0 and 12.7 min for B-LC. These time lags give an effective diffusion coefficient of hydrogen of 5.81 $\pm$ 1.02 $\times$ 10-10 m2.s-1 and 9.74 $\pm$ 2.37 $\times$ 10-10 m2.s-1 for LC and B-LC, respectively. These effective diffusion coefficients are one order higher than the ones reported in the literature (4.5 $\times$ 10-11 m2.s-1 [89] and 7.4 $\times$ 10-11 m2.s-1 [39]) because of the charging conditions in this present work (high current density and acidic solution with NH4SCN). It results in a higher activity of hydrogen and traps are faster filled compared to the corresponding observations reported in previous studies [89, 39]. Additionally, since boron has a repulsive interaction with hydrogen, fewer traps need to be filled in B-LC, which explains the difference in the effective diffusion coefficient of hydrogen between LC and B-LC. Then, the number of possible traps can be determined using equation (10), and the results are presented in table 2.
@@
Figure 9: hermodynamic equilibrium hydrogen fraction in the different crystalline defects as a function of the temperature for total hydrogen concentration of 10 appm (dotted lines), 100 appm (dashed lines), and 1000 appm (solid lines) in a) LC and b) B-LC. The blue lines correspond to the fraction of hydrogen from peak 1 (mostly in the lattice), the red lines correspond to the fraction of hydrogen into the trapping sites of peak 2, the green lines correspond to the fraction of hydrogen into trapping sites of peak 3 and the orange lines correspond to the fraction of hydrogen in trapping sites of peak 4. Figures c and d are zooms of figures a and b for low hydrogen fraction, respectively. Hydrogen distribution at different trapping sites for different total hydrogen concentrations for e) LC and f) B-LC.
Figs. 9.a-d show that all traps are not filled at thermodynamic equilibrium when the total hydrogen concentration is below 1000 appm for the two steels. For comparison, the total hydrogen concentration determined from the peak integration and presented in C is below 700 appm. For the LC steel, hydrogen segregates into trap sites of peak 4 because the trapping energy is lower compared to the other trapping sites. Trap sites of peak 4 are mostly located where boron is segregating (i.e.: PAGBs). This observation also explains why this type of GB is the weakest microstructure feature in hydrogen environments for steels [27, 26]. In addition, even if the permeation experiments revealed fewer possible trap sites ($N_{P_{2}}$ and $N_{P_{3}}$) for B-LC compared to LC, the fraction of hydrogen in trap sites 2 and 3 are higher in B-LC than in LC. Consequently, similar concentrations of hydrogen in traps from peaks 2 and 3 are obtained in the two steels at thermodynamic equilibrium
@@ 6 Conclusion
In this work, the effect of boron segregation on the trapping behavior of hydrogen has been investigated in steel both experimentally and theoretically. First, the microstructure of boron-doped (B-LC) and boron-free (LC) steels has been studied using EBSD, synchrotron X-ray measurements, and correlative TKD-APT measurements to observe any change due to the presence of boron. While the dislocation density, grain size, and grain boundary distributions are similar for LC and B-LC, a pronounced segregation of boron in PAGBs is observed for the B-LC steel. When both steels are pre-charged with hydrogen and measured through TDS measurements, fewer peaks are detected for B-LC. From this behavior, the peak corresponding to hydrogen being trapped in PAGBs can be distinguished from hydrogen trapped in the other microstructural defects. Ab initio calculations demonstrate a strong attraction between boron and GBs (interaction energy of -2.38 eV for a $\Sigma 5(210)[001]$ GB structure), stronger than the interaction between hydrogen and GBs (interaction energy of -0.46 eV for the same GB structure). Additionally, when the GB contains boron solute, the interaction between hydrogen and this defect becomes repulsive. Finally, the equilibrium hydrogen concentration in different defects has been approximated using the McLean formalism with the trapping energy deduced from the TDS measurements and the effective hydrogen diffusion coefficients obtained from the permeation tests. The latter experiment has shown that hydrogen has a higher diffusion coefficient in B-LC compared to LC because fewer traps can be filled by hydrogen in B-LC. Different hydrogen partitioning is observed depending on temperature and total hydrogen concentration. We observed that the concentration of hydrogen in the lattice will be increased for B-LC compared to LC at thermodynamic equilibrium because boron influences the fraction of hydrogen in trap sites pertaining to PAGBs. This effect can induce different failure mechanisms described in different models proposed to explain hydrogen embrittlement. In a situation where traps of the lowest trapping energies like PAGBs are passivated by the presence of boron, it can modify the material’s resistance against hydrogen embrittlement, which is a planned future investigation.
@@ Appendix A EBSD and ECCI mapping of LC and B-LC
The measurements of the average grain size of LC and B-LC have been performed with the EBSD maps from 10.a and 10.b, respectively. For a better representation, a larger surface has been analyzed (surface of at least 2.7 ×$\times 10^{5}$), and a step size of 0.4 $\mu$m has been used for the acquisition of both maps. For the ECCI, the scanning electron microscope (SEM) was operated with an acceleration voltage of 30 kV and a probe current of 5 nA. We inverted the signal from the backscattered electron detector to obtain an image contrast similar to bright-field conditions for transmission electron microscopy images [90].

### v3

@@ 3.1 Experimental details
The incorporation of hydrogen into $LC$ and $B-LC$ has been performed electrochemically, using a three-electrode system on 10$\times$10$\times$1 mm3 samples. The procedure consisted of imposing a current of -25 mA.cm-2 for 1 h in a 0.1 M $H_{2}SO_{4}$ aqueous solution with 0.3 wtof NH4SCN at room temperature. After charging, the oxide layers were removed from the sample using 600 SiC grinding paper before the TDS measurements. For all conditions, the time between the end of the electrochemical charging and the beginning of the temperature ramping is lower than 5 min, which is performed in an infrared furnace from 298 K to 1173 K with a mass spectrometer using the G8 GALILEO apparatus from Bruker®. The heating rate $\varphi$ varied between 8 $K.min^{-1}$ and 32 $K.min^{-1}$ to locate where hydrogen is trapped in the microstructure.
@@ 3.2 Results
Figure 4: TDS spectra of LC and B-LC heated for different temperature ramping at a) 8 K.min-1, b) 12 K.min-1, c) 16 K.min-1, d) 20 K.min-1, e) 24 K.min-1, and f) 32 K.min-1. The sum of all deconvoluted curves is represented for all spectra to compare the difference between the fit and the experiment.
@@
Fig. 5.a presents the Kissinger plot of the different peaks shown in fig. 4 for LC and B-LC steels. The first peak from the TDS spectra of fig. 4 should correspond to mobile hydrogen in the lattice (from the slope of the blue curve). The energy obtained is -0.16 eV and is lower than the theoretically expected activation barrier for diffusion of hydrogen in martensitic steel or $\alpha$-iron, which is between -0.04 eV and -0.1 eV [71, 42, 69]. The main reason is that the determination of this energy comes from the temperature at which the hydrogen desorption rate $\phi$ is the highest. When the heating rate ($\varphi$) is low, this temperature should come around or below room temperature [47], the starting temperature from our experiments. The second peak (red curves) has been attributed to hydrogen trapped at GBs in previous work [41, 47]. Our results show a fourth peak which pertains to hydrogen desorbing from PAGBs. Consequently, this second peak should mostly be related to the desorption of hydrogen trapped in grain boundaries from martensite transformation, which follows the KS OR, and has the same parent austenite grain. The trapping energy deduced from our experiment is -0.26 eV for hydrogen segregating into this type of trapping site (deduced from the slope of the red curve). This energy is in the range of the ones reported in the work of Galindo et al. [43], but outside the range given in Frappart et al. [42, 68] (fig. 5.b). The slope of the green curve gives the average trapping energy of -0.34 eV, and should represent mostly hydrogen in dislocations according to the literature [41, 47]. This value is in the range of the energies given in previous studies [42, 48, 43]. The last peak, observed only for the boron-free steel (orange curve in fig. 5.b) should be hydrogen in PAGBs with a trapping energy of -0.37 eV. This value is close to one reported in the work of Bhadeshia [48] but higher than the one reported in the work of Frappart et al. [42].
@@
In addition, hydrogen could also be trapped in the 2% of retained austenite observed in LC and B-LC steels and in or at carbides [48, 43, 52], precipitates observed by ECCI and synchrotron X-ray measurements. When hydrogen is trapped in austenite, a peak from the TDS spectra above 573 K should be obtained [52], which is not the case, neither for LC nor for B-LC. Different possible hypotheses could explain this behavior and are presented further. The austenite, which is too compressed to transform into martensite, is becoming less attractive site for hydrogen to segregate, even at the interface with the matrix its concentration becomes too low. Such effect is observed in the case of edge dislocations for BCC and FCC metals, where the stress field of the compressed area induces a negative binding energy for hydrogen, and favors its segregation into the tensile area of the defect [72, 73]. Another possible explanation is that the thin layer of this phase reduced the temperature range for hydrogen desorption from it and the small fraction makes it impossible to observe an apparent desorption peak by TDS measurements.
A similar observation is deduced for carbide: its concentration is too low to observe an apparent peak by TDS measurements. A deconvolution of the TDS spectra considering these defects could be done near peaks 2 and 3, but might lead to an inaccurate overfit. Another peak has been found for hydrogen in microvoids in previous works, which has an amplitude significantly lower than the other peaks [41, 47]. However, no additional peak has been observed in either of the two steels. One possible explanation is that hydrogen recombines in these microvoids, making it impossible for hydrogen to desorb at the temperature range studied in this work [47].
@@ 4.1 Computational details
While hydrogen can segregate into any crystalline defects, including PAGBs, boron is only in high concentration in PAGBs. Therefore, the difference in behavior on the hydrogen desorption rate described with TDS measurements should be due to the interaction between hydrogen and boron at this crystalline defect. Consequently, ab initio calculations have been performed to understand this interaction by determining the interaction energies between solutes and the structural defect. have been carried out using the Vienna Ab initio Simulation Package (VASP) code [74, 75, 76] in the generalized gradient approximation (GGA) using the PBE functional [77]. A plane-wave cut-off of 500 eV has been taken for all calculations. The convergence tolerance of the atomic force is 0.01 eV/Å and the Brillouin-zone integration was made using Methfessel–Paxton smearing [78]. The magnetic state of $\alpha$-Fe is ferromagnetic and a $4\times 4\times 4$ supercell containing 128 Fe atoms has been used for the bulk calculations. For the reciprocal space integration, a $\Gamma$-centered Monkhorst-Pack [79] k-point sampling of $8\times 8\times 8$ has been used for bulk $\alpha$-Fe calculations. The lattice constant and magnetic moments of the ferromagnetic $\alpha$-Fe were computed as 2.834 Å and 2.2 $\mu_{\rm B}$/atom, respectively, in agreement with the previous studies [80]. The supercell containing the $\alpha$-$\Sigma$5(210)[001] grain boundary consisted of 80 Fe-atoms. For the reciprocal space integration, a $\Gamma$-centered Monkhorst-Pack [79] k-point sampling of $5\times 10\times 16$ has been used to ensure convergence. The solute atoms H, C, and B have firstly been incorporated in different locations of the defect (in interstitial sites $I_{1}$ and $I_{2}$), represented in fig. 6. Then, additional calculations have been performed to describe the interaction between a boron-doped GB with an additional solute atom. In this situation, boron has been incorporated in $I_{1}$ and a second solute atom (hydrogen and boron) has been incorporated in interstitial site $I_{2}$ or in interstitial sites in the vicinity of the GB (in octahedral and tetrahedral sites).
@@ 4.2 Results
The interaction energy between solutes and boron-doped GB ($E^{\rm inter}_{\rm X-B-GB}$) is presented in fig. 7.b. In this configuration, additional boron is attracted to the GB with an interaction energy lower than -0.4 eV in $O_{2}$. This interaction energy suggests that boron can be easily accumulated in GBs, which facilitates the formation of borides when the nucleation barrier is low enough [32, 33]. In contrast, hydrogen is repelled from the GB in this configuration when inserted in interstitial sites. This result agrees with previous calculations, arguing that boron improves grain boundary cohesion and hydrogen exhibits an ion-like character in metal, including at grain boundary in $\alpha$-iron [37, 38]. For these calculations, when hydrogen atoms are initially inserted in octahedral sites, they are moving to their closest stable tetrahedral site. When a second boron atom is initially inserted in the tetrahedral site, it moves towards its closest stable octahedral site.
@@ 5 Hydrogen concentration in steels at thermodynamic equilibrium
The hydrogen permeation study has been conducted with a Kelvin probe-based detection method developed in-house at room temperature [86, 87]. On the side of the sample designated for hydrogen detection, a dry nitrogen atmosphere is maintained, and a 100 nm Pd coating was applied using a Leybold Univex 450 physical vapor deposition apparatus. The hydrogen-entry side of the sample has been subjected to the same galvanostatic condition as samples for TDS measurements. Time lags of 11.6 min, 15.1 and 17.6 min are obtained for LC and 6.8 min, 8.0 and 12.7 min for B-LC. These time lags give an effective diffusion coefficient of hydrogen of 5.81 $\pm$ 1.02 $\times$ 10-10 m2.s-1 and 9.74 $\pm$ 2.37 $\times$ 10-10 m2.s-1 for LC and B-LC, respectively. These effective diffusion coefficients are one order higher than the ones reported in the literature (4.5 $\times$ 10-11 m2.s-1 [88] and 7.4 $\times$ 10-11 m2.s-1 [42]) because of the charging conditions in this present work (high current density and acidic solution with NH4SCN). It results in a higher activity of hydrogen and traps are faster filled compared to the corresponding observations reported in previous studies [88, 42]. Additionally, since boron has a repulsive interaction with hydrogen, fewer traps need to be filled in B-LC, which explains the difference in the effective diffusion coefficient of hydrogen between LC and B-LC. Then, the number of possible traps can be determined using equation (8), and the results are presented in table 2.
@@
Figure 9: Thermodynamic equilibrium hydrogen fraction in the different crystalline defects as a function of the temperature for total hydrogen concentration of 10 appm (dotted lines), 100 appm (dashed lines), and 1000 appm (solid lines) in a) LC and b) B-LC. The blue lines correspond to the fraction of hydrogen from peak 1 (mostly in the lattice), the red lines correspond to the fraction of hydrogen into the trapping sites of peak 2, the green lines correspond to the fraction of hydrogen into trapping sites of peak 3 and the orange lines correspond to the fraction of hydrogen in trapping sites of peak 4. Figures c and d are zooms of figures a and b for low hydrogen fraction, respectively. Hydrogen distribution at different trapping sites for different total hydrogen concentrations for e) LC and f) B-LC.
Figs. 9.a-d show that all traps are not filled at thermodynamic equilibrium when the total hydrogen concentration is below 1000 appm for the two steels. For the LC steel, hydrogen segregates into trap sites of peak 4 because the trapping energy is lower compared to the other trapping sites. Trap sites of peak 4 are mostly located where boron is segregating (i.e.: PAGBs). This observation also explains why this type of GB is the weakest microstructure feature in hydrogen environments for steels [29, 30]. In addition, even if the permeation experiments revealed fewer possible trap sites ($N_{P_{2}}$ and $N_{P_{3}}$) for B-LC compared to LC, the fraction of hydrogen in trap sites 2 and 3 are higher in B-LC than in LC. Consequently, similar concentrations of hydrogen in traps from peaks 2 and 3 are obtained in the two steels at thermodynamic equilibrium
@@ 6 Conclusion
In this work, the effect of boron segregation on the trapping behavior of hydrogen has been investigated in steel both experimentally and theoretically. First, the microstructure of boron-doped (B-LC) and boron-free (LC) steels has been studied using EBSD, synchrotron X-ray measurements, and correlative TKD-APT measurements to observe any change due to the presence of boron. While the dislocation density, grain size, and grain boundary distributions are similar for LC and B-LC, pronounced segregation of boron in PAGBs is observed for the B-LC steel.
Then, both steels have been pre-charged with hydrogen and measured through TDS measurements. The disappearance of one peak has been observed for B-LC compared to LC, suggesting it corresponds to hydrogen trapped in PAGBs and can accordingly be distinguished from hydrogen trapped in the other microstructural defects.
Ab initio calculations demonstrate a strong attraction between boron and GBs (interaction energy of -2.38 eV for a $\Sigma 5(210)$ GB structure), stronger than the interaction between hydrogen and GBs (interaction energy of -0.46 eV for the same GB structure). Additionally, when the GB contains boron atoms, the interaction between hydrogen and this defect becomes repulsive.
Finally, the equilibrium hydrogen concentration in different defects has been approximated using the McLean formalism with the trapping energy deduced from the TDS measurements and the effective hydrogen diffusion coefficients obtained from the permeation tests. The latter experiment has shown that hydrogen has a higher diffusion coefficient in B-LC compared to LC because fewer traps can be filled by hydrogen in B-LC. Different hydrogen partitioning is observed depending on temperature and total hydrogen concentration. Compared to LC, the concentration of hydrogen in the lattice will be increased for B-LC at thermodynamic equilibrium because boron saturates the fraction of hydrogen in trap sites pertaining to PAGBs. This effect can induce different failure mechanisms described in different models proposed to explain hydrogen embrittlement. In situations where strong traps like PAGBs for hydrogen are passivated by boron, the material’s resistance against hydrogen embrittlement can be modified: this hypothesis is a planned future investigation.
@@ Appendix A EBSD and ECCI mapping of LC and B-LC
The measurements of the average grain size of LC and B-LC have been performed with the EBSD maps from 10.a and 10.b, respectively. For a better representation, a larger surface has been analyzed (surface of at least 2.7 ×$\times 10^{5}$ $\mu^{2}$), and a step size of 0.4 $\mu$m has been used for the acquisition of both maps. For the ECCI, the scanning electron microscope (SEM) was operated with an acceleration voltage of 30 kV and a probe current of 5 nA. We inverted the signal from the backscattered electron detector to obtain an image contrast similar to bright-field conditions for transmission electron microscopy images [89].